THE VALUE OF x, WHEN  2x+4.3x+1=288

Grade:9

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
From the Given data 2x+4.3x+1=288
Asked to find the value of x
So 2x+12x+1=288
14x+1=288
14x=288-1=287
14x=287
x=(287/14)
x=41/2
So x=20.5